OpenStudy (anonymous):

Guys, seriously help, what is 2+2

OpenStudy (cggurumanjunath):

seriously it is 999

OpenStudy (anonymous):

2+2 = (1+1) +(1+1) = 4

OpenStudy (anonymous):

it is 4, who doasnt know that?????????? okay i am pissed off i had never heard such a question...

OpenStudy (anonymous):

am i being too rude? ok then sorry but, if u want... USE A CALCULATOR!!!!!!!

OpenStudy (anonymous):

@faizah is this crazy or not?????

OpenStudy (anonymous):

omg!...are u in elementary?... 2+2=4 :o

OpenStudy (anonymous):

@i.am.a.troll this is dumb

OpenStudy (anonymous):

i know right???!!?!?!?!?!

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

@i.am.a.troll ur name certainly proves everything :p

OpenStudy (anonymous):

yep, you truly are a troll!